# On Space-Filling Input Design for Nonlinear Dynamic Model Learning: A Gaussian Process Approach

### v1

@@ IV-B Space-filling Promoting Cost Function
The space-filling promoting cost function is constructed based on the dataset $\mathcal{D}_{N}=\{z_{j}\}_{j=1}^{N}$ generated by the parameterized input signal $u(k;\theta)$ after applying it to the assumed model of the system (1). Based on the dataset $\mathcal{D}_{N}$, one has following data matrix $Z=\mathrm{vec}(z_{1},\ldots,z_{N})\in\mathbb{R}^{N\times n_{z}}$ where $z_{i}=\mathrm{vec}(x(i),u(i;\theta))$, $i\in\mathbb{I}_{i=1}^{N}$.
@@
Given a region of interest $\tilde{\mathcal{Z}}\subseteq\mathcal{Z}$, the anchor dataset is characterized by $\mathcal{D}_{\mathrm{I}}:=\{\tilde{z}_{i}=\mathrm{vec}(\tilde{x}_{i},\tilde{u}_{i})\}_{i=1}^{M}$, which contains $M$ pre-defined anchor points from $\mathcal{Z}$ to capture the dynamics, thereby gridding the region of interest $\tilde{\mathcal{Z}}$.
We place a GP prior on the latent hypothetical model $\hat{x}(k+1)=\hat{f}(x(k),u(k))$ of the known data-generating system (1), which results in $\hat{f}\sim\mathcal{GP}\left({\mu}({z}),{\kappa}\left({z},{z}^{\prime}\right)\right)$. The kernel function $\kappa\left({z},{z}^{\prime}\right)$ satisfies the following assumption:
@@
The kernel function $\kappa(\|z-z^{\prime}\|)=\kappa(z,z^{\prime})$ is chosen as a degenerating, monotonically decreasing function w.r.t. $\|z-z^{\prime}\|$.
The decreasing monotonicity of $\kappa(z,z^{\prime})$ indicates that the association between the anchor points and $Z$ diminishes with the increase of the Euclidean distance $\|z-z^{\prime}\|$. The kernel is usually chosen from the exponential family, e.g., the Squared Exponential Automatic Relevance Determination (SEARD) $\kappa\left({{z}},{{z}}^{\prime}\right)\!=\!\sigma_{\mathrm{f}}^{2}\exp\left(-\frac{1}{2}({{z}}\!-\!{{z}}^{\prime})^{\top}{\Lambda}^{-1}({{z}}\!-\!{{z}}^{\prime})\right)\!$ is taken as a prior with $\sigma_{\mathrm{f}}^{2}\in\mathbb{R}^{+}$ and $\Lambda=\mathrm{diag}(\ell_{1}^{2},...,\ell_{n_{\mathrm{z}}}^{2})\in\mathbb{R}^{n_{\mathrm{z}}\times n_{\mathrm{z}}}$.
@@
We would like to find a parametrized input signal $u(k;\theta)$ such that the resulting dataset $\mathcal{D}_{N}(\theta)$ is expected to achieve the $1/\epsilon$ density, ensuring a good coverage within the region of interest $\tilde{\mathcal{Z}}$. To this end, when taking the dynamical system (1) into account, the posterior variance at each anchor point can be rewritten in a $\theta$-dependent form:
@@
$\quad\mathcal{D}_{N}(\theta)\!:=\!\{{Z}^{(j)}_{\theta}\!=\!\mathrm{vec}(x(j),u(j;\theta))\}_{j=1}^{N},$ (13d)
where
@@
The implementation of the proposed method is summarized in Algorithm 1.
@@
  Step 1: Select the number of samples $N$, learning rate $\alpha$ and learning threshold $\delta$. Set the iteration index $\iota=0$. Initialize $u(k;\theta^{(0)})$ with $\theta^{(0)}=\theta_{0}$;
@@
     Apply $u(k;\theta^{(\iota)})$ on (1) and obtain $\mathcal{D}_{N}(\theta^{(\iota)})$;
@@
  Step 5: Set $\hat{\theta}\leftarrow\theta^{(\iota)}$, and adopt $u(k;\hat{\theta})$ as the space-filling input signal.
@@ IV-C Theoretical Analysis
  • The anchor dataset $\mathcal{D}_{\mathrm{I}}$ has the $1/\epsilon$ density;
@@
The first step is to prove that the posterior variance at $\tilde{z}_{i}$: $\mathrm{Var}(\hat{f}\!\mid\!\tilde{z}_{i},\!\mathcal{D}_{N})\!=\!{\kappa}(\tilde{z}_{i},\tilde{z}_{i})\!-\!{\kappa}(\tilde{z}_{i},\!Z)K_{N}^{-1}{\kappa}(\tilde{z}_{i},Z)^{\top}\!=\!0$ if and only if there exists $j\in\mathbb{I}_{1}^{N}$ such that $\tilde{z}_{i}=z^{(j)}$.
@@
(Necessity) We show the necessity part by contradiction. Suppose that $\tilde{z}_{i}\notin\mathcal{D}_{N}$, and ${\kappa}(\tilde{z}_{i},\tilde{z}_{i})-{\kappa}(\tilde{z}_{i},Z)K_{N}^{-1}{\kappa}(\tilde{z}_{i},Z)^{\top}=0$ still holds. Since the kernel function $\kappa(\cdot,\cdot)$ is positive definite, the augmented matrix $\left[\begin{smallmatrix}K_{N}&\kappa(\tilde{z}_{i},Z)^{\top}\\ \kappa(\tilde{z}_{i},Z)&\kappa(\tilde{z}_{i},\tilde{z}_{i})\end{smallmatrix}\right]\succ 0$ holds for all $\tilde{z}_{i}\notin\mathcal{D}_{N}$. By the Schur complement condition for positive definiteness, the Schur complement of $K_{N}$ must be positive, that is, ${\kappa}(\tilde{z}_{i},\tilde{z}_{i})-{\kappa}(\tilde{z}_{i},Z)K_{N}^{-1}{\kappa}(\tilde{z}_{i},Z)^{\top}>0$, $\forall\tilde{z}_{i}\notin\mathcal{D}_{N}$. This contradicts with the assumption, which implies that only if $\tilde{z}_{i}\in\mathcal{D}_{N}$, $\mathrm{Var}(\hat{f}\mid\tilde{z}_{i},\mathcal{D}_{N})=0$ holds.
@@
(Sufficiency) Assume that $\forall i\in\mathbb{I}_{1}^{M}$, the anchor point $\tilde{z}_{i}$ has at least one data $z^{(j)}$ on it. Based on the proof in the first step, we have $\mathrm{Var}(\hat{f}\mid\tilde{z}_{i},\mathcal{D}_{N})=0$. This leads to $\mathcal{V}(\mathcal{D}_{N})=0$
@@
Based on Lemma 1, subsequently, we show the $\epsilon$-space filling property of a dataset $\mathcal{D}_{N}(\theta)$ which is generated by the optimized input applied to the dynamical system. To achieve this, the following condition is assumed to be satisfied for a set of parametrized input signal $\mathscr{U}(\theta)$ given in (2):
@@
There exists a sufficiently flexible parametrization for $\mathscr{U}(\theta)$ such that it can realize any possible input sequence, for example, $u(k;\theta)=\theta_{k},\forall k\in\mathbb{Z}$.
@@
Consider system (1) which satisfies Condition 1 and the parameterized input signal $u(k;{\theta})$ which satisfies Conditions 2 and 4. The GP-based space-filling promoting cost function is defined as (14) on the basis of the $\epsilon$-space filling anchor dataset $\mathcal{D}_{\mathrm{I}}=\{\tilde{z}_{i}\}_{i=1}^{M}$, with a kernel function $\kappa(z,z^{\prime})$ satisfying Condition 3. Then the dataset $\mathcal{D}_{N}(\hat{\theta})$ generated by the optimized input signal $u(k;\hat{\theta})$ achieves the $\epsilon$-space filling property in the region of interest $\tilde{\mathcal{Z}}$ with a finite data length $N\geq M\times(T_{\mathrm{d}}+1)$, where $\hat{\theta}=\mathrm{arg\,min}_{\theta\in\Theta}~{}\mathcal{W}(\theta;\mathcal{D}_{N}(\theta)).$
@@
On the basis of Condition 1, it is obvious that starting from any $z_{j}=\mathrm{vec}(x(j),u(j))\in\mathcal{Z}$, it takes 1 step to reach $x(j+1)$, and another $T_{\mathrm{d}}$ steps to reach any anchor point $\tilde{z}_{i}=\mathrm{vec}(\tilde{x}(i),\tilde{u}(i))\in\tilde{\mathcal{Z}}$, $\forall i\in\mathbb{I}_{1}^{M},j\in\mathbb{I}_{1}^{N}$. Hence, all $M$ anchor points can be explored with a finite data length $N\geq M\times(T_{\mathrm{d}}+1)$. Condition 4 guarantees that the parametrization of $\mathscr{U}({\theta})$ is sufficiently flexible to generate any value of the input. Combining again with Condition 1, it can be deduced that there exists a parameter vector $\theta\in\Theta$ such that the corresponding input signal $u(k;{\theta})$ can generate a dataset $\mathcal{D}_{N}(\theta)$ that satisfies $\mathcal{V}(\mathcal{D}_{N})=0$. Then according to Lemma 1, one can readily conclude that the dataset $\mathcal{D}_{N}(\hat{\theta})$ generated by $u(k;\hat{\theta})$ optimized based on the cost function (14) is guaranteed to have the $1/\epsilon$ density w.r.t. $\tilde{\mathcal{Z}}$. ∎
@@
Theorem 1 implies that minimizing (14) leads to the $\epsilon$-space filling regardless of the hyperparameter choice. It is worth mentioning that our proposed space-filling input design strategy is not concerned with the GP mean estimate, as the standard GP regression does. In practice, the selection of $\sigma_{\mathrm{f}}^{2}$ and $\Lambda$ can make the problem easier to optimize. Moreover, in the finite data case, the length scale $\Lambda$ should be chosen considering both the interval between anchor points and the size of $N$ to ensure sufficient coverage.
@@ V-A Linear system example
$\dot{x}=Ax+Bu$ (16)
where $A=\begin{bmatrix}0&1\\ -0.3&-0.5\end{bmatrix}$ and $B=\begin{bmatrix}0\\ 1\end{bmatrix}$. The model (16) is discretized using zero-order hold on the input with a sample time of $T_{s}=1$s. The region of interest is selected as $\tilde{\mathcal{Z}}=[-2,2]\times[-2,2]$. The input sequence is a fully parameterized signal $u(k;\theta)=\theta_{k}$ with $N=40$ samples, where the input value at each time step is treated as an optimization variable. Initially, $\theta_{k}^{(0)}\sim\mathcal{N}(0,1)$ for all $k\in\mathbb{I}_{1}^{N}$. Figure 1 illustrates the space-filling performance under different numbers of anchor points when the optimized cost tends to 0 with $Q$ set as an identity matrix. It can be observed that $\mathcal{D}_{N}(\theta)$ is lower than the theoretically guaranteed space-filling property ($\epsilon=2.8,1.4,0.94$) for 20 Monte Carlo realizations. Moreover, as $M$ increases, the radius of the largest ball decreases, indicating a corresponding increase in density, thereby leading to improved space-filling performance.

### v2

@@ IV-B Space-filling Promoting Cost Function
The proposed space-filling promoting cost function is constructed based on the dataset $\mathcal{D}_{N}\!=\!\{z_{j}\}_{j\!=\!1}^{N}$ generated by the parameterized input signal $u(k,\theta)$ after applying it to the known model of the system (1). Based on the dataset $\mathcal{D}_{N}$, one has following data matrix $Z\!=\!\mathrm{vec}(z_{1},\ldots,z_{N})\in\mathbb{R}^{N\times n_{z}}$ where $z_{i}\!=\!\mathrm{vec}(x(i),u(i,\theta))$, $i\in\mathbb{I}_{i\!=\!1}^{N}$.
@@
Given a region of interest $\tilde{\mathcal{Z}}\subseteq\mathcal{Z}$, the anchor dataset is characterized by $\mathcal{D}_{\mathrm{I}}\!:=\!\{\tilde{z}_{i}\!=\!\mathrm{vec}(\tilde{x}_{i},\tilde{u}_{i})\}_{i\!=\!1}^{M}$, which contains $M$ pre-defined anchor points from $\mathcal{Z}$ gridding the region of interest $\tilde{\mathcal{Z}}$.
We place a GP prior on the latent hypothetical model $\hat{x}(k+1)\!=\!\hat{f}(x(k),u(k))$ of the known data-generating system (1) which results in $\hat{f}_{\nu}\sim\mathcal{GP}\left({\mu}_{\nu}({z}),{\kappa}\left({z},{z}^{\prime}\right)\right),\nu\in\mathbb{I}_{1}^{n_{\mathrm{x}}}$, where all $\hat{f}_{\nu}$ share the same kernel ${\kappa}\left({z},{z}^{\prime}\right)$ and input space. This constitutes to $\hat{f}\!=\!\mathrm{vec}(\hat{f}_{1},...\hat{f}_{n_{\mathrm{x}}})$. The kernel function $\kappa\left({z},{z}^{\prime}\right)$ satisfies the following assumption:
@@
The kernel function $\kappa(z,z^{\prime})\!=\!\kappa(\|z-z^{\prime}\|)$ is chosen to be a positive definite kernel that is monotonically decreasing w.r.t. $\|z-z^{\prime}\|$.
The decreasing monotonicity of $\kappa(z,z^{\prime})$ indicates that the association between the anchor points and $Z$ diminishes with the increase of the Euclidean distance $\|z-z^{\prime}\|$. The kernel is usually chosen from the exponential family, e.g., the Squared Exponential Automatic Relevance Determination (SEARD) $\kappa\left({{z}},{{z}}^{\prime}\right)\!=\!\sigma_{\mathrm{f}}^{2}\exp\left(-\frac{1}{2}({{z}}\!-\!{{z}}^{\prime})^{\top}{\Lambda}^{-1}({{z}}\!-\!{{z}}^{\prime})\right)\!$ is taken as a prior with $\sigma_{\mathrm{f}}^{2}\in\mathbb{R}^{+}$ and $\Lambda\!=\!\mathrm{diag}(\ell_{1}^{2},...,\ell_{n_{\mathrm{z}}}^{2})\in\mathbb{R}^{n_{\mathrm{z}}\times n_{\mathrm{z}}}$, where both $\sigma_{\mathrm{f}}^{2}$ and $\Lambda$ are treated as hyperparameters.
As all $\hat{f}_{\nu}$ share the same kernel and input space, they yield identical posterior variance $\forall\nu\in\mathbb{I}_{1}^{n_{\mathrm{x}}}$. For simplicity, we drop the dimension indexing and refer to $\mathrm{Var}(\hat{f}\mid\tilde{z}_{i},\mathcal{D}_{N})$ instead of $\mathrm{Var}(\hat{f}_{\nu}\mid\tilde{z}_{i},\mathcal{D}_{N})$ for the remainder of the paper.
@@
We would like to find a parametrized input signal $u(k,\theta)$ such that the resulting dataset $\mathcal{D}_{N}(\theta)$ is expected to achieve the $1/\epsilon$ density, ensuring a good coverage within the region of interest $\tilde{\mathcal{Z}}$. To this end, when taking the dynamical system (1) into account, the posterior variance at each anchor point can be rewritten in a $\theta$-dependent form:
@@
$\quad\mathcal{D}_{N}(\theta)\!:=\!\{z_{j}\!=\!\mathrm{vec}(x(j),u(j,\theta))\}_{j=1}^{N},$ (12d)
with $Z_{\theta}\!=\!\mathrm{vec}(z_{1},...,z_{N})$, where
@@
The implementation of the proposed method is summarized in Algorithm 1. The computational complexity per optimization step is dominated by the posterior variance calculation which scales as $\mathcal{O}(N^{3}+N^{2}(n_{\mathrm{x}}+n_{\mathrm{u}}+M))$, and is primarily determined by the input length $N$. As the proposed method is performed offline, the computational cost remains acceptable even for moderately large $N$ and system dimensions.
Remark 3
Compared to recent works [15, 16], the proposed GP-based cost does not rely on matching a predefined distribution. Matching a specific distribution can be restrictive as the desire to cover a space does not impose a specific distribution over that space. Hence, leaving room to stay closer to the natural preferences of the system dynamics by not specifying such a target distribution is beneficial to reach the space covering goal. Furthermore, the proposed GP-based cost remains smooth and differentiable, enabling gradient-based optimization under arbitrary input parameterizations.
@@
  Step 1: Select the number of samples $N$, step size $\alpha$ and threshold $\delta$. Set the iteration index $\iota\!=\!0$. Initialize $u(k,\theta^{(0)})$ with $\theta^{(0)}\!=\!\theta_{0}$;
@@
     Apply $u(k,\theta^{(\iota)})$ on (1) and obtain $\mathcal{D}_{N}(\theta^{(\iota)})$;
@@
  Step 5: Set $\hat{\theta}\leftarrow\theta^{(\iota)}$, and adopt $u(k,\hat{\theta})$ as the space-filling input signal.
@@ IV-C Theoretical Analysis
  • The anchor dataset $\mathcal{D}_{\mathrm{I}}$ has $1/\epsilon$ density;
@@
First we prove that the posterior variance at $\tilde{z}_{i}$: $\mathrm{Var}(\hat{f}\mid\tilde{z}_{i},\!\mathcal{D}_{N})\!=\!{\kappa}(\tilde{z}_{i},\tilde{z}_{i})\!-\!{\kappa}(\tilde{z}_{i},\!Z)K_{N}^{-1}{\kappa}(\tilde{z}_{i},Z)^{\top}\!=\!0$ if and only if there exists a $j\in\mathbb{I}_{1}^{N}$ such that $\tilde{z}_{i}\!=\!z^{(j)}$.
@@
(Necessity) We show the necessity part by contradiction. Suppose that $\tilde{z}_{i}\notin\mathcal{D}_{N}$, and ${\kappa}(\tilde{z}_{i},\tilde{z}_{i})-{\kappa}(\tilde{z}_{i},Z)K_{N}^{-1}{\kappa}(\tilde{z}_{i},Z)^{\top}\!=\!0$ still holds. Since the kernel function $\kappa(\cdot,\cdot)$ is positive definite, the augmented matrix $\left[\begin{smallmatrix}K_{N}&\kappa(\tilde{z}_{i},Z)^{\top}\\ \kappa(\tilde{z}_{i},Z)&\kappa(\tilde{z}_{i},\tilde{z}_{i})\end{smallmatrix}\right]\succ 0$ for all $\tilde{z}_{i}\notin\mathcal{D}_{N}$. As $K_{N}$ is positive definite, by the Schur complement, ${\kappa}(\tilde{z}_{i},\tilde{z}_{i})-{\kappa}(\tilde{z}_{i},Z)K_{N}^{-1}{\kappa}(\tilde{z}_{i},Z)^{\top}>0$, $\forall\tilde{z}_{i}\notin\mathcal{D}_{N}$. This contradicts with the assumption, hence implying that only if $\tilde{z}_{i}\in\mathcal{D}_{N}$, $\mathrm{Var}(\hat{f}\mid\tilde{z}_{i},\mathcal{D}_{N})\!=\!0$ holds.
@@
(Sufficiency) Assume that $\forall i\in\mathbb{I}_{1}^{M}$, the anchor point $\tilde{z}_{i}$ has at least one data $z^{(j)}$ on it. Based on the proof in the first step, we have $\mathrm{Var}(\hat{f}\mid\tilde{z}_{i},\mathcal{D}_{N})\!=\!0$. This leads to $\mathcal{V}(\mathcal{D}_{N})\!=\!0$.
@@
Based on Lemma 1, subsequently, we show the $\epsilon$-space filling property of a dataset $\mathcal{D}_{N}(\theta)$ which is generated by the optimized input applied to the dynamical system. To achieve this, the following condition is assumed to be satisfied for a set of parametrized input signals $\mathscr{U}(\Theta)$ given in (2):
@@
There exists a sufficiently flexible parametrization for $\mathscr{U}(\Theta)$ such that for each time step $k$, the input value $u(k)$ can be freely specified without constraints by appropriately choosing $\theta$, for example, $u(k,\theta)\!=\!\theta_{k},\forall k\in\mathbb{Z}$.
@@
Consider system (1) which satisfies Condition 1 and the parameterized input signal $u(k,{\theta})$ which satisfies Conditions 2 and 4. The GP-based space-filling promoting cost function is defined as (13) on the basis of the $\epsilon$-space filling anchor dataset $\mathcal{D}_{\mathrm{I}}\!=\!\{\tilde{z}_{i}\}_{i\!=\!1}^{M}$, with a kernel function $\kappa(z,z^{\prime})$ satisfying Condition 3. Then, the dataset $\mathcal{D}_{N}(\hat{\theta})$ generated by the optimized input signal $u(k,\hat{\theta})$ achieves the $\epsilon$-space filling property in the region of interest $\tilde{\mathcal{Z}}$ with a finite data length $N\geq M\times(T_{\mathrm{d}}+1)$, where $\hat{\theta}\!=\!\mathrm{arg\,min}_{\theta\in\Theta}~{}\mathcal{W}(\theta;\mathcal{D}_{N}(\theta)).$
@@
On the basis of Condition 1, it is obvious that starting from any $z_{j}\!=\!\mathrm{vec}(x(j),u(j))\in\mathcal{Z}$, it takes 1 step to reach $x(j+1)$, and another $T_{\mathrm{d}}$ steps to reach any anchor point $\tilde{z}_{i}\!=\!\mathrm{vec}(\tilde{x}(i),\tilde{u}(i))\in\tilde{\mathcal{Z}}$, $\forall i\in\mathbb{I}_{1}^{M},j\in\mathbb{I}_{1}^{N}$. Hence, all $M$ anchor points can be explored with a finite data length $N\geq M\times(T_{\mathrm{d}}+1)$. Condition 4 guarantees that the parametrization of $\mathscr{U}({\Theta})$ is sufficiently flexible to generate any value of the input. Combining again with Condition 1, it can be deduced that there exists a parameter vector $\theta\in\Theta$ such that the corresponding input signal $u(k,{\theta})$ can generate a dataset $\mathcal{D}_{N}(\theta)$ that satisfies $\mathcal{V}(\mathcal{D}_{N})\!=\!0$. Then according to Lemma 1, one can readily conclude that the dataset $\mathcal{D}_{N}(\hat{\theta})$ generated by $u(k,\hat{\theta})$ optimized based on the cost function (13) is guaranteed to have the $1/\epsilon$ density w.r.t. $\tilde{\mathcal{Z}}$. ∎
@@
Theorem 1 implies that minimizing (13) leads to the $\epsilon$-space filling regardless of the hyperparameter choice. It is worth mentioning that our proposed space-filling input design strategy is not concerned with the GP mean estimate, as the standard GP regression does. In practice, the selection of $\sigma_{\mathrm{f}}^{2}$ and $\Lambda$ can make the problem easier to optimize. Moreover, in the finite data case, the length scale $\Lambda$ should be chosen considering both the interval between anchor points and the size $N$ to ensure sufficient coverage.
@@ V-A Linear system example
$\dot{x}=Ax+Bu,\quad A\!=\!\begin{bmatrix}0&1\\ -0.3&-0.5\end{bmatrix},\quad B\!=\!\begin{bmatrix}0\\ 1\end{bmatrix}$ (15)
Model (15) is discretized using zero-order hold on the input with a sample time of $T_{s}\!=\!1$s. The region of interest is selected as $\tilde{\mathcal{Z}}\!=\![-2,2]\times[-2,2]$. The input sequence is a fully parameterized signal $u(k,\theta)\!=\!\theta_{k}$ with $N\!=\!40$ samples, where the input value at each time step is treated as an optimization variable. Initially, $\theta_{k}^{(0)}\sim\mathcal{N}(0,1)$ for all $k\in\mathbb{I}_{1}^{N}$. Figure 1 illustrates the space-filling performance under different numbers of anchor points when the optimized cost tends to 0 with $Q$ set as an identity matrix. It can be observed that $\mathcal{D}_{N}(\theta)$ achieves a lower $\rho$ than the theoretically guaranteed space-filling bound ($\epsilon\!=\!2.8,1.4,0.94$) over the 20 Monte Carlo realizations. Moreover, as $M$ increases, the radius of the largest ball decreases, indicating a corresponding increase in density, thereby leading to improved space-filling performance.